Assume France and Libya can both produce grain and dates, and that the only limited resource is the farming labor force, meaning that land, water, and all other resources are plentiful in both countries. Each farmer in France can produce 10 t of grain or 5 t of dates in a season. Each farmer in Libya can also produce 10 t of grain or 25 t of dates. Please answer the four questions. Which country has the absolute advantage in producing dates? Which country has the absolute advantage in producing grain? O France O Libya O Neither O France O Neither O Libya
Which country has the comparative advantage in producing dates? Which country has the comparative advantage in producing grain? O France O Libya O Neither O Neither O Libya O France
Answer #1

1) Absolute Advantage:

Absolute advantage is the ability of a country to produce a good or service at a lower per unit cost as compared to any other country that produces same good or service.

Libya is able to produce 25 t of dates as compared to France which is able to produce only 5 t of dates. So, Libya has an absolute advantage in the production of dates.

2) Neither country has an absolute advantage in the production of grains because both is able to produce 10 t of grains.

3) Comparative Advantage

A country has a comparative advantage in producing that good if the opportunity cost of producing that good is lower in that country as compared to another country.

France;

5t of dates = 10t of grains

1t of date = 10/5 = 2t of grains

1t of grain = 5/10 = 0.5t of date

Opportunity cost of 1 t of date is 2 t of grains and 1t of grain is 0.5t of date.

Libya;

25t of dates = 10t of grains

1t of date = 10/25 = 0.4t of grain

1t of grain = 25/10 = 2.5t of date

Opportunity cost of 1 t of date is 0.4 t of grains and 1t of grain is 2.5t of date.

Libya has comparative advantage in the production of date because of lower opportunity cost of Libya in the production of dates as compared to France.

4) France has comparative advantage in the production of grain because of lower opportunity cost of France in the production of grain as compared to Libya.